OpenStudy (anonymous):

If the bases are 1/2 and 1/4, and the height is 1, what is the area? 1/8 sq. units 3/8 sq. units 3/4 sq. units

OpenStudy (campbell_st):

I'm sorry, they way the question is written, it makes no sense. Perhaps you need to post an image of the question or include the type of shape you are working on

OpenStudy (anonymous):



OpenStudy (anonymous):

does this help any?

OpenStudy (anonymous):

so then 1/2 +1/4 = .75

OpenStudy (anonymous):

divide that by 2 and you will have your area

OpenStudy (campbell_st):

the formula is \[A = \frac{1}{2} \times ( b_{1} + b_{2}) \times h\] great so leave it as a fraction 3/4 now find 1/2 of 3/4 or 1/2 x 3/4 = leave the answer as a fraction

OpenStudy (campbell_st):

when you get the answer, to finish multiply by the height h = 1 so Area = 1 x 1/2 x 3/4

OpenStudy (anonymous):

is that 3/8 then?

OpenStudy (anonymous):

yes the answer is 3/8

OpenStudy (anonymous):

thank you so much